OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 4(x - 6) = -2(x - 3)

OpenStudy (anonymous):

I think Hannapults is the best name ever. Anyways onto the math!! We can distribute the 4 and -2. Our new equation: 4x-24=-2x-6 Does this make sense so far?

OpenStudy (anonymous):

yes i think so, and thankyou, i like you name too

OpenStudy (anonymous):

*your

OpenStudy (anonymous):

:D Anyways, its the rule that you can only have and x on one side of the equation. Because it's easier to add I'll add 2x to the right side of our equation HOWEVER since it's equal you'll need to add 2 to the other side too. so our equation is now 4x+2x-24=-2x-6+2x Solving this gives us 6x-24=-6 Still understanding?

OpenStudy (anonymous):

yep, you make it sound so easy, lol xD

OpenStudy (anonymous):

xD It'll get easy. It took me a week to understand this. Anyways, now that we have only a number on the right side we can remove our -24 from the left side. But it's still even, so we have to add it to the right side too. 6x-24+24=-6+24 Our new equation is 6x=18 Can you do the last step or is it still unclear?

OpenStudy (anonymous):

I can do the last step :D thankyou so much for helping me :)

OpenStudy (anonymous):

No problem! Good luck with the rest of your work!!

OpenStudy (anonymous):

ok, X=3 right? just making sure